8x+7y=-8
-3x+y=32
solve by sub method
ty for your help

Answer by Maths68(1474)


8x+7y=-8.........(1)
-3x+y=32..........(2)
By substitution method:-
Take (2)
-3x+y=32
Subtract y from both sides of the equation.
-3x+y-y=32-y
-3x=32-y
Divide by -3 both sides of the equation
-3x/-3=(32-y)/-3
x=(32-y)/-3
Now substitute the value of x in equation (1)
8x+7y=-8
8[(32-y)/-3]+7y=-8
(256-8y)/-3+7y=-8
Multiply both sides of the equation by -3
256-8y-21y=24
256-29y=29
subtract 256 from both sides of the equation
256-256-29y=24-256
-29y=-232
Divide by -29 both sides of the equation
-29y/-29=-232/-29
y=232/29
y=8
Put the value of y in any of two equations (1) or (2)
8x+7y=-8
8x+7(8)=-8
8x+56=-8
Subtract 56 from both sides of above equation
8x+56-56=-8-56
8x=-64
Divide by 8 both sides of the above equation
8x/8=-64/8
x=-8
Solution set (x,y)=(-8,8)
Check:-
Put the values of x and y in any two equations (1) or (2)
-3x+y=32
-3(-8)+8=32
24+8=32
32=32
Now
8x+7y=-8
8(-8)+7(8)=-8
-64+56=-8
-8=-8